A consumer purchases two goods, food and clothing. The utility function is U=FC where



F denotes the amount of food consumed and C the amount of clothing. The price of food



is PF=2,000 Birr, the price of clothing is PC=4,000 Birr, and the income M=80,000. Find



the equilibrium of this consumer.

1
Expert's answer
2022-01-04T10:20:26-0500

In equilibrium MUf/Pf = MUc/Pc and 2,000F + 4,000C = 80,000.

C/2,000 = F/4,000,

F = 2C,

2,000×2C + 4,000C = 80,000,

8,000C = 80,000,

C = 10 units,

F = 20 units.